# Hyperellipsoid Density Sampling: Exploitative Sequences to Accelerate High-Dimensional Optimization

### v1

@@ Abstract.
The method was evaluated against Sobol, a standard QMC method, using differential evolution (DE) on the 29 CEC2017 benchmark test functions. The results show statistically significant improvements in solution geometric mean error ($p<0.05$), with average performance gains ranging from $3\%$ in $30D$ to $37\%$ in $10D$. This paper demonstrates the efficacy of HDS as a robust alternative to QMC sampling for high-dimensional optimization.
@@ 2.2.12. Transform to Hyperellipsoid
  • $\mathbf{C}_{k}$ is the cluster centroid vector, $\mathbf{c}_{k}$, added (broadcast) to every row of the scaled and rotated matrix.
@@ 3.2. Computational Efficiency
$n\_{\text{samples}}$ Avg Time (s) Std Dev (s) $\text{CI}_{95}$
@@
SamplesSamples $D$ $Ratio_{time}$ $p$-Value
@@ 4. Ease of Use & Implementation
HDS is implemented as a streamlined Python package, following the conventions of Scipy’s QMC module. It uses a small set of standard library imports for minimal external dependency. The open-source code is available at [github.com/jgsoltes/Hyperellipsoid-Density-Sampling] for full use and reproducibility.

### v3

@@ Abstract.
The method was evaluated against Sobol, a standard QMC method, using differential evolution (DE) on the 29 CEC2017 benchmark test functions. The results show statistically significant improvements in solution geometric mean error ($p<0.05$), with average performance gains ranging from $3\%$ in $30D$ to $37\%$ in $10D$. This paper demonstrates the efficacy of HDS as a robust alternative to uniform QMC sampling for high-dimensional optimization.
@@ 2.2.12. Transform to Hyperellipsoid
  • $\mathbf{C}_{k}$ is the cluster centroid vector, $\mathbf{c}_{k}$, added to every row of the scaled and rotated matrix.
@@ 3.2. Computational Efficiency
NN Avg Time (s) Std Dev (s) $\text{CI}_{95}$
@@
NN $D$ $Ratio_{time}$ $p$-Value
@@ 4. Ease of Use & Implementation
HDS is implemented as a streamlined Python package, following the conventions of Scipy’s QMC module. It uses a small set of standard library imports for minimal external dependency. The open-source code is available at [github.com/jgsoltes/hdim-opt] for full use and reproducibility.